Can someone help me on this problem?
I need to solve and check the following equation:
12x - 6 = 11x
Thanks, Ashley

Answer by checkley71(8403)

12X-6=11X OR 12X-11X=6 OR X=6
PROOF 12*6-6=11*6 OR 72-6=66 OR 66=66